### v2

# Identifying and Addressing Delusions for Target-Directed Decision Making

Abstract
We are interested in target-directed agents, which produce targets during decision-time planning, to guide their behaviors and achieve better generalization during evaluation. Improper training of these agents can result in delusions: the agent may come to hold false beliefs about the targets, which cannot be properly rejected, leading to unwanted behaviors and damaging out-of-distribution generalization. We identify different types of delusions by using intuitive examples in carefully controlled environments, and investigate their causes. We demonstrate how delusions can be addressed for agents trained by hindsight relabeling, a mainstream approach in for training target-directed RL agents. We validate empirically the effectiveness of the proposed solutions in correcting delusional behaviors and improving out-of-distribution generalization.
1 Introduction
One important use of Reinforcement Learning (RL) is to train agents with skills generalizable to target tasks for application, after learning from limited training tasks, where the evaluation and training task distributions may differ (Ada et al., 2024). Despite numerous efforts to bridge the “generalization gap” over discrepancies between the two sets of tasks, generalization abilities of RL agents remain largely unsatisfactory. This means that learned skills become less useful to tasks in practice, greatly limiting the scope of applications and research interests. Recent works attributed this to existing agents’ lack of reasoning abilities to face Out-Of-Distribution (OOD) changes (Di Langosco et al., 2022), as most agents either operate solely based on intuition (including model-free methods and offline model-based methods) (Kahneman, 2017). These suggestions led to the development of agents that reason adaptively in novel situations.
Thus, some decision-time planning agents are developed, which make use of their planning outcomes, which we call targets, to adapt in OOD scenarios (Alver & Precup, 2022). Specifically, targets are self-generated (sub-)goals produced during decision-time planning. The “target-directed” agents make use of generators to propose targets, as well as optional estimators that evaluate the favorability of the targets, to assist solving the given tasks. These processes however provide no inherent assurance to the selected targets.
In psychiatry, delusions depict the inability to reject false beliefs (Kiran & Chaudhury, 2009), which in the RL context translates to false beliefs that are natural results of an agent’s learning. Particularly problematic for target-directed agents, when false beliefs cannot be resolved through training, delusional behaviors may appear in the form of chasing nonexistent or unreachable targets, or those beyond safety constraints (Bengio et al., 2024). Delusions limit the agents’ situational understanding, and lead to even worse generalization in the target tasks. Delusions stem from the incoordination of the generator and the estimator, respectively: problematic targets should not be generated, and even if they were, they should not be favored(Corlett, 2019). While delusions have various sources, this paper focuses on the misevaluation by the estimator caused by improper training processes, a combination of inadequate update rules and inappropriate training data distributions.
In this paper, we demonstrate how to preemptively address failure modes induced by delusions in the target tasks by training the estimators properly. With clear diagnoses in the designed environments, we identify types of delusions, and investigate more proper designs of the training processes. We then applied our idea to hindsight relabeling and validated in experiments that our strategies led to better understanding of the patterns of delusions and significant reduction in delusional behaviors, resulting in significant improvements in OOD generalization performance.
2 Preliminaries
RL & Problem Setting.: An RL agent takes actions to interact with an environment to gain cumulative reward. The interaction may be modeled as a Markov Decision Process (MDP) $\mathcal{M}\equiv\langle\mathcal{S},\mathcal{A},P,R,d,\gamma\rangle$, where $\mathcal{S}$ and $\mathcal{A}$ are the set of states and actions, $P:\mathcal{S}\times\mathcal{A}\to\text{Dist}(\mathcal{S})$ is the state transition function, $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}$ is the reward function, $d:\mathcal{S}\to\text{Dist}(\mathcal{S})$ is the initial state distribution, and $\gamma\in(0,1]$ is the discount factor. An agent needs to learn a policy $\pi:\mathcal{S}\to\text{Dist}(\mathcal{A})$ that maximizes the value, i.e. the expected discounted cumulative reward $\mathbb{E}_{\pi,P}[\sum_{t=0}^{T_{\perp}}\gamma^{t}R(S_{t},A_{t},S_{t+1})|S_{0}\sim d]$, where $T_{\perp}$ denotes the time step when the episode terminates. Often, environments are partially observable, which means, instead of a state, the agent receives an observation $x_{t+1}$, with which the agent needs to infer the state from the history.
Figure 1: Target-Directed Frameworks: A generator proposes candidate target(s). An estimator can be used to evaluate and select the favorable among candidates.
Target-Directed RL offers a perspective to better identify delusions. It abstracts existing methods that belong to the intersection of decision-time planning and generalized goal-conditioned RL (Nasiriany et al., 2019; Zadem et al., 2024). The framework emphasizes an algorithmic design, where a generator proposes candidate targets, and an (optional) estimator evaluates them and select one target for (optionally target-conditioned) policies to execute (Dayan & Hinton, 1992; Parr & Russell, 1997). The two components correspond to the belief formation and belief evaluation systems, respectively, whose incoordination causes delusions in the human brain (Kiran & Chaudhury, 2009), as shown in Fig. 1.
Hindsight Relabeling is a mainstay approach to train conditioned policies, estimators, as well as temporally-extended generators. Training target-directed agents with only contemporary targets (the ones being followed) can lead to poor performance (Dai et al., 2021), especially for agents that propose their own targets, where contemporary targets may be low-quality, hard to achieve, or lack in diversity (Moro et al., 2022; Davchev et al., 2021). Hindsight Experience Replay (HER) addresses this by creating a buffer of training data for the learner to sample from and filling it with relabeled transitions, which are augmented with alternative targets (Andrychowicz et al., 2017). HER augments a transition $\langle x_{t},a_{t},r_{t+1},x_{t+1}\rangle$ with an additional observation $x^{\odot}$ (or its some encoding), the relabeled target. Each (re)labeled transition constitutes a source-target pair, with one decision point for the current step and another for the relabeled target. Relabeling strategies, which correspond to the sources of $x^{\odot}$, are critical for the performance of HER-trained agents (Shams & Fevens, 2022). Most popular choices are trajectory-level, meaning $x^{\odot}$ comes from the same trajectory as $x_{t}$ and are thus conducted at the end of each episode. These include “future”, where $x^{\odot}=x_{t^{\prime}}$ with $t^{\prime}>t$, and “episode”, with $0\leq t^{\prime}\leq T_{\perp}$.
SSM In most works, failure modes of target-directed agents are often overlooked, possibly due to the complexity of benchmark environments. To better identify the causes, and provide intuitive examples, we craft a set of fully-observable environments based on the MiniGrid-BabyAI framework (Chevalier-Boisvert et al., 2018b; Hui et al., 2020), named SwordShieldMonster (SSM for short). In SSM, the agent moves one step at a time in four directions to navigate fields with episode-terminating, randomly placed lava traps. The density of traps is controlled by a difficulty parameter $\delta$, which guarantees a viable path to success, i.e. acquiring a sparse terminal reward. Agents must collect a sword and a shield, both randomly placed, before reaching the “monster” guarding the treasure. Reaching the monster prematurely ends the episode without reward. Visualizations of SSM are in Fig. 2, along with examples of delusional behaviors discussed later. The sword and shield are acquired by moving to their respective grid cells, and agents cannot revert the process. This introduces temporary unreachability in the state structure, as non-terminal states are not fully traversable from one another. Semantically, this segments SSM states into four equivalence classes, formed by two binary statuses: the agent’s possession of the sword and shield. For example, $\langle 0,1\rangle$ denotes sword not acquired, shield acquired.”
3 Delusions in Target-Directed RL
Delusions can lead to agent chasing outcomes that damage the agents’ generalization in the target tasks or those with catastrophic consequences. We would like to identify delusions, develop corresponding mitigating strategies, s.t. through learning on the training tasks, the agents can identify and avoid delusional behaviors in the target tasks. We look into the two high-level components of target-directed frameworks, i.e. the generator and the estimator, for the sources of delusions and the possible causes.
(a) E.3.2.1 (current state in $\langle 1,1\rangle$)
(b) E.3.2.2 (current state in $\langle 1,1\rangle$, target in $\langle 0,0\rangle$)
Figure 2: Visualizations of Delusional Behaviors in SSM: The agent is marked as a red triangle, lava traps as orange squares, and monsters as green loops. In both cases, the estimators, lacking exposure to problematic targets during training, failed to estimate their reachability. This led to the delusion that these problematic targets (yellow dots), seemingly close to every other state, can be used as proxies to reach the distant task goal quickly.
3.1 Generator Delusions
Most learned generators are subjected to unwanted generalization, and hence will inevitably generate problematic targets, a behavior sometimes referred to as hallucination (Jafferjee et al., 2020). While hallucination can be reduced, they are generally unavoidable. In nature, hallucinations are symptoms of delusions, as the issue of unwanted generalization is unlikely to be resolved by the agent autonomously. While delusional behaviors are not all from problematic targets, generator hallucinations are a major source of risks related to delusions.
Generators are often implemented as neural networks and trained in various ways, where direct causes of problematic target generation could be complicated and hard to enumerate. Hence, we focus on identifying the types of problematic targets based on their characteristics.
3.1.1 Type G.3.1 - Nonexistent
The first type of problematic targets includes those that do not correspond to valid states in the task MDP. They can be classified into: 1) Invalid: semantically invalid targets for the task, such as a target singleton of an imagined SSM observation without the agent position; 2) Impossible: semantically valid but unreachable targets like a SSM target cell surrounded by lava traps, (such as in Fig. 2 a). G.3.1 targets only cause delusional behavior if the estimator produces matching Type E.3.2.1 delusions, to be discussed later. Imperfectly generated targets corresponding to real states may not be problematic, as estimators may filter out imperfections, possibly using a state encoder.
3.1.2 Type G.3.1.1 - Temporarily Unreachable
The second type includes those that correspond to states in the task MDP, but (some) cannot be reached from the current state. Unlike G.3.1, these targets are only temporarily inappropriate. There are some notable subtypes, such as: 1) Irreversible: targets once reachable but are now blocked due to a past transition. For instance, after acquiring the sword in SSM, the agent transitions from class $\langle 0,0\rangle$ to $\langle 1,0\rangle$, sealing off access to $\langle 0,0\rangle$ or $\langle 0,1\rangle$; 2) Segregated: targets that can only be reached from specific initial states that are not available to the agent. For example, an agent spawned in class $\langle 1,0\rangle$ cannot reach states in $\langle 0,0\rangle$ or $\langle 0,1\rangle$. Naïvely applying relabeling strategies, e.g.episode”, to generator training, can be a major cause of G.3.1.1 hallucinations. Similarly, G.3.1.1 may not result in delusional behaviors if the matching Type E.3.2.2 delusions are not produced by the estimator.
G.3.1.1 targets may appear more frequently if we do not intentionally train the generators to avoid generating them. For example, a conditional target generator which learns from “episode” will be more likely to produce G.3.1.1 targets, such as the one visualized in Fig. 2 b). This intuition applies to training methods beyond HER as well. Importantly, note that G.3.1.1 targets do not exist in all MDPs.
Problematic targets such as G.3.1 and G.3.1.1 can cause delusional behaviors of chasing unreachable targets, if they are favored during decision-time.
3.2 Estimator Delusions
An estimator evaluates the favorability of proposed targets, based on some relationship between the current state and the candidates, which often involve assessing the implementability of targets, represented as, e.g., distances, etc. Acting as a firewall, the estimator of a target-directed agent can assist the selection among the proposed targets, bearing the potential to filter out delusional behaviors. Agents without estimators, e.g. Hafner et al. (2022), must accept proposed targets unconditionally, and thus are at significant risk of delusions (chasing G.3.1 and G.3.1.1). Importantly, the incorporation of estimators merely delegates the responsibility of addressing delusional behaviors to the estimator.
Estimators’ delusional beliefs about targets can directly cause delusional behaviors. For an estimator to learn to reject delusions, we need 1) effective learning rules that can punish the favorability of problematic targets, and 2) training data that can counteract delusions through the learning rules. A lack of either can result in blind spots of knowledge about targets, delusional or non-delusional.
Update Rules: we expect estimators to learn to reject problematic, undesirable or unsafe targets, by evaluating them with low favorability. However, such expectation may not be achieved, if the estimators’ update rules do not have the key characteristic that punishes unachievable / unsafe targets, when these targets are sampled during training for the estimator to learn from.
Training Data Distribution Even with capable update rules, delusions can arise from training data distributions that lack the exposure to the data counteracting false beliefs in the estimator. For example, some training data distributions suffer from low-diversity, e.g., training target-directed agents relabeling mechanisms (Dai et al., 2021). But relabeling isn’t necessarily the solution: 1) Certain relabeling strategies intrinsically cause delusions. For instance, “future” only relabels future observations, and thus only exposes a learner to reachable targets, confusing the estimator when a “past” target is generated at decision time; 2) Trajectory-level relabeling is also problematic. Short trajectories, common in many training procedures, prevent estimators from learning to evaluate distant targets. Procedures like “episode” limit exposure, risking delusions. Short trajectories can be a product of experimental design (initial state distributions, episode lengths) or innate environmental characteristics (density of terminal states). For example, to train useful skills in a hard-exploration problem, a popular technique is to relax the initial state distribution of the training tasks, thus producing shorter but more efficient trajectories (Erraqabi et al., 2021); In some other cases, with dense terminal states or short Maximum Episode Lengths (MELs), each experienced trajectory generally only covers a limited portion of the state space.
With these root causes, we identify types of estimator delusions based on the targets they evaluate. We use the identifiers E.3.2, E.3.2.1, and E.3.2.2 for their resulting delusional behaviors.
3.2.1 Type E.3.2 - Misevaluating Non-Delusional Targets
The standalone type of delusions includes false estimations about non-delusional targets. E.3.2 delusions can lead to undesirable or delusional targets being favored. E.3.2(-induced) behaviors can hurt the agents’ generalization abilities.
3.2.2 Type E.3.2.1 - Misevaluating G.3.1 Targets
This type of delusion appears when an estimator mis-evaluates the favorability of a G.3.1 target. If effective learning rules are present, this can still be caused by the lack of training data sampling procedure (e.g., existence of relabeled transitions with G.3.1 targets in HER). In Fig. 2 a), an example is visualized. Note that the resulting E.3.2.1 behaviors, i.e. those chasing selected G.3.1 targets, can be potentially catastrophic if the G.3.1 targets are beyond safety constraints.
3.2.3 Type E.3.2.2 - Misevaluating G.3.1.1 Targets
Similar to E.3.2.1, in Fig. 2 b), an example of delusional behavior is visualized, a result of a combination of G.3.1.1 (temporarily unreachable) and E.3.2.2. E.3.2.2 behaviors can hurt the agents’ generalization abilities.
4 Addressing Delusions: Strategies Applied to Hindsight Relabeling
Identifying delusions enables mitigation strategies. We propose general solutions for existing target-directed agents, coming from a spectrum of training procedures. Given the popularity of HER-based training and the generality behind hindsight relabeling, we provide examples of applying our strategies to address delusions in HER-based training, demonstrating empirical benefits.
Assume we are dealing with a hypothetical target-directed agent with a dual-component architecture as in Fig. 1, learning both components exclusively from hindsight-relabeled transitions. The generator proposes targets corresponding to potentially distant states, and the estimator learns the favorability of targets (via an oracle or more complex methods) by punishing the unachieved or unsafe targets present in the training data. To address delusions, we combine update rules and training data distribution. By assuming this framework, we skip discussing update rules, as they depend on the agent’s design.
4.1 Assistive Strategies for Addressing Delusions
We first introduce two strategies to address delusions, which are then materialized as two hindsight relabeling strategies for HER. Later, we compare the ideas behind the new strategies with existing ones, and study how the two groups can be combined to create mixtures for estimators and hybrids for the generator-estimator duo.
The following strategies seek to expand the support of the training data distribution, to include those source-target combinations involving G.3.1 and G.3.1.1 targets, respectively. With the assumed update rules, this extra exposure is expected to address E.3.2.1 and E.3.2.2 estimator delusions.
4.1.1 generate”: Let Estimators Learn About Candidates (to be generated)
The first strategy, named “generate”, is to let the estimator learn about targets that could be proposed at decision time, s.t. it could figure out preemptively that problematic targets are not favorable.
In Zhao et al. (2024), the authors identified a delusional behavior resulting from, E.3.2.1 delusions in the language of this paper, and they proposed to train the estimator additionally with candidate targets proposed by the generator. With HER, we can transform this auxiliary loss into a Just-In-Time (JIT) HER strategy that, when a transition is sampled for training, relabels it with a conditionally generated target by the generator. “generate” is particularly effective for training estimators, as the estimators will get exposed to all kinds of problematic targets (that the generator could offer [^1] When a generator learns to generate mostly viable targets, “generate” will be primarily effective at dealing with the E.3.2 and E.3.2.1 (G.3.1s are unlikely to be eradicated). “generate” requires the use of the generator, thus it incurs additional computational burden.
Adapting “generate” beyond HER should be straightforward. Whenever a target is involved (during estimator learning), we can replace it with generated targets.
4.1.2 pertask”: Let Estimators Learn About Experienced Targets
The second strategy, namely “pertask”, is to expose the estimator to targets that were experienced before, s.t. it could figure out that some targets are not favorable from the current state.
We materialize “pertask” into an assistive atomic relabeling strategy that relabels transitions with observations from the same training task, sampled across the entire memory. “pertask” works on a task level, creating more diverse source-target pairs than trajectory-level strategies like “future” & “episode”, which bias the training data distribution toward shorter-distance source-target pairs. This approach additionally brings exposure to the estimator to learn against E.3.2.2 delusions and to the long-distance source-target pairs, thus mitigating E.3.2 caused by short trajectories.
It is worth noting that “pertask” also biases the training data distribution, making the agent spread out its efforts into learning the source-target pairs that are potentially far away from each other. Despite the increment to training data diversity, long-distance pairs are less likely to contribute to better decision-making compared to the shorter-distance in-episode ones offered by “episode”, as shown later in experiments.
Adapting “pertask” beyond HER requires algorithmic designs recording all past observations.
Both strategies help the estimator figure out the features shared by problematic targets, s.t. OOD delusions can also be identified. Note that “pertask” cannot be used to address E.3.2.1 delusions. However, “generate” can be used for E.3.2.2 if the generator generates the respective problematic targets.
4.2 Mixtures
Creating a mixture of sources of training data increases the diversity of source-target combinations. For HER specifically, each atomic strategy, enumerated in Tab. 4.2, exhibits a tradeoff in estimation accuracy among all sorts of source-target pairs, including short-distance and long-distance ones involving only valid targets, and those involving problematic targets.
A mixture of more-than-one atomic strategies in certain proportions while relabeling (Nasiriany et al., 2019; Yang et al., 2021a), can be used to achieve a tradeoff in HER-based training, s.t. the shortcomings of each atomic strategy are mitigated by the introduction of others.
{NiceTabular}
c|m[l]0.25|m[l]0.5|m[l]0.25 Strategies Advantages   Disadvantages   Gist
episode” Efficient for estimator to learn close-proximity relationships Can cause G.3.1.1 targets when used to train generators; When used exclusively to train estimator, 1) cannot handle E.3.2.2; 2) prone to E.3.2 - cannot learn well from short trajectories Creates training data with source-target pairs sampled from the same episodes
future” Can be used to learn a conditional generator with temporal abstractions In addition to the shortcomings of “episode” (those for estimators only), this additionally causes E.3.2 when used as the exclusive strategy for estimator training Creates training data with temporally-ordered source-target pairs from the same episodes
generate” Addresses E.3.2.1 with data diversity (also E.3.2.2 when generator produces G.3.1.1) Relies on the generator with additional computational costs; Potentially low efficiency in learning non-delusional relationships. Augments training data to include candidate targets proposed at decision time
pertask” Addresses estimator delusions (E.3.2.2 & E.3.2 for long-distance pairs) Can cause extensive G.3.1.1 targets if used to train generators; low efficiency in learning close-proximity source-target relationships Augments training data to include targets that were experienced
Table 1: Hindsight Relabeling Strategies & The Ideas Behind: “episode” and “future” are widely used as they increase sample efficiency in non-delusional cases significantly; “generate” and “pertask”, proposed in this paper, are effective against delusions, useful in specific scenarios.
When the training budget is fixed, i.e., training frequency, batch sizes, etc., stay unchanged, the mixing proportions of strategies pose a tradeoff to the investment towards different kinds of source-target pairs, and the resulting accuracies. In experiments, we show that, assisting “episode” with “generate” and “pertask” often results in better performance in estimator training, striking a balance between the investment of training budgets in non-delusional and delusional estimations.
Mixtures of multiple relabeling strategies can also be extended to other training procedures, and can also be implemented by independent training procedures based on different sources.
4.3 Hybrid Strategies - A $2$-Slotted Approach
While an estimator needs the exposure to problematic targets to counteract delusions, a generator benefits from learning to generate useful targets, and should avoid exposure to problematic targets. Thus, generators and estimators often have conflicting needs during training.
Each source of training data has its own biases. Atomic hindsight relabeling strategies in HER, as shown in Tab. 4.2, can simultaneously help one component while hindering another. Instead of trying to tradeoff the needs of both components with a single source of training data, we propose a hybrid ($2$-slotted) approach, allowing the generator and estimator to receive data tailored to their needs, through two independent relabeling processes. This combination of $2$ slots and mixtures can produce various hybrid strategies, as demonstrated in Sec. 6.4.
Separating training data for the generator and estimator based on their needs is straightforward in training procedures beyond HER.
In the Appendix, we provide summaries for the design of mitigation strategies on agents beyond HER.
5 Related Works
Target-Directed Agents. Dual-component frameworks are explicitly investigated with temporal abstraction for purposes such as path planning (Nasiriany et al., 2019), OOD generalization (Zhao et al., 2024) and task decomposition (Nair & Finn, 2020; Zadem et al., 2024). Davchev et al. (2021) employs the similar framework to assist exploration, without using HER for training.
Hindsight Relabeling (Andrychowicz et al., 2017) is crucial for enhancing sample efficiency in target-directed RL, as it enables learning from failed experiences (Dai et al., 2021). From the perspective of source-target pair distribution, from which we similarly developed this paper, improvements were proposed, including a divergence maximization approach (Zhang & Stadie, 2022), addressing distributional change (Bai et al., 2023), and multistep relabeling strategies (Yang et al., 2021b), etc. Kuang et al. (2020) proposed to prioritize rarely-seen achieved transitions by utilizing the density distribution. HER’s success was centered on its performance gains via sample efficiency (in non-delusional cases), around which most follow-up works revolved as well.
While, performance deficiencies of target-directed frameworks can be resulted from multiple factors, which include delusions. Shams & Fevens (2022) studied the performance of atomic strategies from the view of sample efficiency, without looking into the failure modes. Deshpande et al. (2018) detailed experimental techniques in sparse reward settings using “future”. In (Yang et al., 2021a), a mixture strategy similar to “generate” improved non-delusional estimations, though the impact on delusions was not explored, possibly because of the lack of an appropriate RL environment. Nasiriany et al. (2019) used a mixture of up to $3$ atomic strategies for producing effective training data in a single-task setting, while its contribution to addressing delusions was not investigated. The performance of existing target-directed agents are often limited by their exclusive reliance on “future” or “episode(He et al., 2020), whose delusions this paper intends to address.
Delusions in RL. Delusions can stem from various other improper designs. For example, Lu et al. (2018) identified delusions caused by the limitations of function approximators for greedy policies in modelfree RL; Often accompanying hallucinations, delusions can arise in non-target-directed frameworks as well. In an offline planning framework such as Dyna, state hallucinations can irreversibly destabilize value estimations (Jafferjee et al., 2020). In Di Langosco et al. (2022); Shah et al. (2022), the authors classified goal misgeneralization, a class of delusional behaviors with which an agent competently pursues an undesired target leading in novel test situations. Zhao et al. (2024) gave first examples of delusional behaviors caused by improper hindsight relabeling.
6 Experiments
To investigate the effectiveness of our strategies against delusions and how they can help target-directed agents achieve better generalization, we choose a combination of environments with clearly defined delusional cases, as well as methods whose estimators can be easily analyzed. This leads to our $4$ sets of experiments, coming from a combination of $2$ environments (one dominantly haunted by G.3.1, and another by G.3.1.1) and $2$ target-directed frameworks. Due to page limit, $3$ out of $4$ sets of experiments are only presented in the Appendix.
6.1 Environments & Settings
For environments, we favor tasks where the identified delusions are present and intuitive to inspect, for which the introduced SSM offers us clear advantage. Introduction and the results on another environment are presented in the Appendix.
For each training seed run, we sample and preserve $50$ training tasks from a distribution of tasks of size $12\times 12$ and difficulty $\delta=0.4$. All agents are trained for $1.5\times 10^{6}$ interactions by randomly selecting one of the $50$ tasks for each training episode. To speed up training, the initial state distributions span all the non-terminal states in each training task. Note that the change of initial states also increases risks of E.3.2.2, due to the presence of dense episode-terminating lava traps and relatively short MELs ($128$ for SSM).
6.2 Evaluative Criteria
We employ the following criteria to investigate the changes in agent’s estimations and behaviors:
Estimation Error - E.3.2, E.3.2.1, E.3.2.2 & Non-Delusional: At each evaluation point, we solve the ground truth distances between states and inspect the agents’ estimation errors. The errors are split based on the source-target pairs, into those involving problematic targets, i.e., E.3.2.1 and E.3.2.2, and those involving only valid targets, which includes the case of E.3.2.
Delusional Behavior Frequencies: We also monitor the frequency of a problematic target (G.3.1 or G.3.1.1) being chosen by the agents, as the result of their delusions in decision-time planning.
Improvement on OOD Generalization Performance: We analyze the changes in agents’ OOD generalization performance, due to the strategies introduced to handle delusions. The evaluation tasks are sampled from a gradient of OOD difficulties - $0.25$, $0.35$, $0.45$ and $0.55$. For aggregated OOD performance, such as in Fig. 3 g), we sample $20$ tasks from each of the $4$ OOD difficulties, and combine the performance across all $80$ episodes, which have a mean difficulty matching the training tasks. By comparing the resulted performances of atomic strategies with those of the newly proposed hybrids, we can also have an intuitive grasp of the “delusion gap”, which is the amount of performance degradation caused by delusions. To maximize difficulty, the initial state is fixed in each novel evaluation task instance: the agents are not only spawned to be at the furthest side of the monster, but also in semantic class $\langle 0,0\rangle$, i.e., without the sword nor the shield in hand.
6.3 Methods
To clearly demonstrate the effectiveness against delusions and the resulting OOD performance boost, we adapted the following target-directed methods:
Skipper: At decision-time, it generates candidate target states that, together with the current state, form a directed graph. The states act as vertices and the edges are pairwise estimations of cumulative rewards and discounts (interchangeable with distances), under its evolving policy. A target state is chosen among candidates after applying value iteration (VI), i.e. the favorability of targets are defined by the $Q$ values of the resulting planned paths. Candidates can be preserved for re-selection of targets (Zhao et al., 2024).
LEAP (in Appendix): Instead of VI, LEAP uses an evolutionary algorithm to evolve sequences of subgoals leading to the task goal, with distances as fitness. The immediate subgoal of the elitest sequence, among all, is chosen to be the target and used to condition a lower-level policy. LEAP is prone to delusions because its planned sequences require all subgoals to be non-delusional (Nasiriany et al., 2019).
For both agents, the generators and estimators are trained with HER, allowing easy switching of relabeling strategies for comparison. To a degree, both agents employ update rules that can lead to rejection of wrong estimations with data exposure (more discussions in Appendix). Targets are acquired by encoding imagined observations from end-to-end learned generators, where problematic propositions can be clearly identified (G.3.1 & G.3.1.1).
6.4 Hindsight Relabeling Strategies
Besides on the basic “episode”, “future” and “pertask”, we conduct comparative studies with the following baselines for fairer comparison (excluding bad generator performance):
  • F-E: “future” for generator, “episode” for estimator;
  • F-P: “future” for generator, “pertask” for estimator;
  • F-G: “future” for generator, with $100\%$ chance using “generate” JIT on estimator;
Additionally, the following hybrid strategies are proposed and compared (more in Appendix):
    F-(E+G): A hybrid strategy to deal with mostly E.3.2.1. “future” for generator, “episode” for estimator with $50\%$ chance using “generate” JIT, resulting in a mixture of $50\%$episode” and $50\%$generate” for estimator;
    F-(E+P): A hybrid strategy for mostly E.3.2.2. “future” for generator, mixture of $50\%$episode” and $50\%$pertask” for estimator;
    F-(E+P+G): A hybrid strategy against both E.3.2.1 and E.3.2.2 behaviors. “future” for generator, mixture of $2/3$episode” and $1/3$pertask” for estimator, with $1/4$ chance using “generate” JIT, resulting in a mixture of $50\%$episode”, $25\%$pertask” and $25\%$generate” for estimator;
(a) G.3.1 Candidate Ratio
(b) E.3.2.1 Estimation Errors
(c) E.3.2.1 Behavior Ratio
(d) Non-Delusional Estim. Errors
(e) G.3.1.1 Candidate Ratio
(f) E.3.2.2 Estimation Errors
(g) E.3.2.2 Behavior Ratio
(h) Aggregated OOD Performance
Figure 3: Skipper on SSM: each mean curve and CI (95%) bar are established over $20$ seed runs. Subfigures b,c,d,f,g,h) share the same legends. a-c) focus on E.3.2.1 delusions: a) Evolving ratio of G.3.1 targets among all candidates at each target selection, throughout training, is presented. Variants are differentiated only with the generator HER strategies; b) Evolving E.3.2.1 delusions measured by $L_{1}$ error in estimated distance throughout training is visualized. The distances to unreachable targets are clipped to be the maximum values that the estimator could output ($16$ in experiments). A more detailed breakdown of this error is in the Appendix; c) The curves represent the frequencies of choosing G.3.1 targets whenever a selection of targets is initiated; d) Final estimation accuracies of non-delusional source-target pairs after training completed, across a spectrum of ground truth distances. In this subfigure, both distances (estimation and ground truth) are conditioned on the snapshot of the final policies. Note that a part of the errors comes from E.3.2; Subfigures e-g) are the E.3.2.2-related counterparts of a-c); h) Each data point represents OOD evaluation performance aggregated over $4\times 20$ newly generated tasks, with mean difficulty matching the training tasks. The decomposed curves can be found in the Appendix.
6.5 Skipper on SSM (Set 1/4)
Generators: For the HER-trained generators, Fig. 3 a), shows that none among “future”, “episode” and “pertask” could directly address G.3.1 target generation with significance. While, Fig. 3 e) indicates that, “future” generates G.3.1.1 delusional targets significantly less frequently than “episode” and “pertask”, as the other two wasted training budget on G.3.1.1 targets, especially “pertask” that brings in even more problematic training samples from long distances. For fairer comparison in the estimators, we only compare variants with “future” for the generator training, i.e., “F-*”.
Estimators: Exclusive use of “episode” for estimators resulted in high accuracy in non-delusional cases (cases involving pairs of non-delusional states, “F-E” in Fig. 3 d)) for source-target pairs of both short and long distances, but low accuracies in E.3.2.1 and E.3.2.2 cases (Fig. 3 b) & f)); Exclusive use of “pertask” results in significantly worse non-delusional short-distance estimates (“F-P” in Fig. 3 d)), yet much better than “episode” (“F-E”) and “future” (“F-F”) in long-distance cases (indicating that trajectory-level “future” and “episode” lead to E.3.2 cases for longer-distance source-target pairs). This is likely because, without the presence of a backbone strategy such as “episode”, the estimator spent too much of its budget learning long-distance non-delusional or E.3.2.2 cases; Similarly, for “generate” (“F-G”), high accuracy is achieved for E.3.2.1 related cases, at the sacrifice of E.3.2.2 and non-delusional cases; All $3$ hybrid non-baseline strategies achieve good estimation accuracies in both short- and long-distance non-delusional estimation, as shown in Fig. 3 d). At the same time, we can observe particularly significant improvement in accuracy in E.3.2.2 delusional estimates in Fig. 3 f), by “F-(E+P)” and “F-(E+P+G)”, the hybrids assisted by “pertask”.
Behaviors: For all compared variants, we can deduce from 3 that generally, less frequent G.3.1.1 generation (e)) and lower E.3.2.2 errors (f)) lead to less frequent delusional behaviors in g), which in turn improves the OOD performance in h); “episode” (“F-E”) showed mediocre evaluation performance (h)) for its good estimation accuracy in short-distance non-delusional cases, but was prone to delusional behaviors; While, despite that “pertask” (“F-P”) showed decent accuracies in longer-distance non-delusional and E.3.2.2 cases (d) & f)), its low accuracy in short-distance non-delusional estimation (d)) devastated the agent to exhibit the lowest performance in h); Similarly, despite that “generate”’s (“F-G”’s) effectiveness in addressing E.3.2.1-induced behaviors (Fig. 3 c)), its resulting bad estimation accuracies in the non-delusional cases destroyed its OOD performance.
All $3$ non-baseline hybrids achieve significantly better performance, as shown in Fig. 3 h). “F-(E+P)” and “F-(E+P+G)”, the hybrids assisted by “pertask”, performed the best in terms of reducing E.3.2.2 delusions (in Fig. 3 f)) and consequently addressed the most delusional behaviors (in Fig. 3 g)). This indicates a major presence of E.3.2.2 behaviors of Skipper’s failure modes on SSM.
6.6 Summary of Experiments & Empirical Guidelines
With the proposed strategies, we saw a reduction in delusional behaviors during training in both Skipper and LEAP (in Appendix), which led to better OOD generalization performance in two sets of environments, posing challenges of G.3.1 and G.3.1.1, respectively (more results in Appendix).
6.7 Empirical Guidelines
Training generators with problematic targets could cause hallucinations, which increases risks for E.3.2.1 and E.3.2.2 induced delusions; Incorporate an estimator to enable rejection of unwanted targets; Use update rules with appropriate properties and try to maximize the diversity of the training data distribution (against E.3.2); Inspect the candidate targets if possible, since their low quality may indicate higher E.3.2.1 risks, which we can mitigate with the idea behind “generate”; Analyze the state structure of the target tasks and identify temporary unreachabilities. These E.3.2.2 risks may indicate the effectiveness of the idea behind “pertask”.
7 Conclusion & Future Work
We discussed the causes and mitigation of delusions, a shared failure mode of target-directed agents. When applying our strategies on HER-trained agents, the resulting hybrid hindsight relabeling approaches flexibly satisfy the needs of the generator and the estimator, while avoiding the naïveté in training processes that could lead to more delusional behaviors.
It is likely that we did not exhaustively identify all potential types of delusions in target-directed agents. We wish to continue the investigation into unidentified causes of failure modes and design more useful skill-learning agents.
8 Reproducibility Statement
The results presented in the experiments are fully-reproducible with the source code published at https://github.com/pwnerharry/delusions.
9 Acknowledgements
Mingde is grateful for the financial support of the FRQ (Fonds de recherche du Québec) and the collaborative efforts during his internship at Borealis AI, Montreal. We appreciate the computational resources allocated to us from Mila and McGill university.
Appendix A Summaries
For better understanding of the contributions of this paper, we provide some brief summaries about the identified delusions / delusional behaviors and the corresponding mitigation strategies. For generator delusions, please check Tab. A. For estimator delusions, please check Tab. A. We hope that these could inspire the application on more target-directed methods.
{NiceTabular}
c|m[l]0.4|m[l]0.4 Types Causes Solutions
G.3.1 The use of neural networks in the generator architecture makes G.3.1 targets almost inevitable Can cause G.3.1.1 targets when used to train generators; No general solution for this type
G.3.1.1 Problematic training procedures that did not isolate the generators from temporarily unreachable targets Isolate generator from learning temporarily unreachable states
Table 2: Causes & Solutions for Generator Delusions
{NiceTabular}
m[l]0.2|m[l]0.4|m[l]0.4 Causes Explanations Solutions
Inappropriate Update Rules Some update rules do not lead to punishment of the favorability of unreachable targets Adjust the learning rules
Bad Training Data Distributions the collected experience was not organized in a way that produced sufficiently diverse source-target distributions to make the agent realize, via appropriate update rules, that wrong beliefs are wrong Adjust how training data distribution is crafted (minding G.3.1 and G.3.1.1).
Table 3: Causes & Solutions for Estimator Delusions
Appendix B Experiments
B.1 Skipper on SSM (Set 1 / 4, cont.)
B.1.1 Breakdown of E.3.2.1 errors
Skipper utilizes policy evaluation based on Temporal Difference updates to learn the cumulative rewards and discounts along the way from one state to another. These recursive update rules based on state-action (Q) values enable the effective learning of estimates conditioned on the evolving policy within the trust regions. However, they are constrained in the sense that they can only learn the unfavorability of hallucinated targets from non-delusional states, not the other way around, if the considerate training procedure is present to provide samples against the delusions. This is particularly problematic for cases involving G.3.1 targets, but not those with G.3.1.1, since there exist source-target pairs in ER where the generated G.3.1.1 targets serve as sources.
Fig. 4 provides a breakdown of the E.3.2.1 errors, into cases depending on if the G.3.1 targets serve as the source or the target in the corresponding relationship estimation on the source-target pairs. From the accuracy gaps between pairs of variants in each case, we can see that, with our proposed hybrid strategies, Skipper can effectively deal with case e), therefore leading to better accuracy in case c) and a).
(a) Overall
(b) from G.3.1 only
(c) to G.3.1 only
(d) G.3.1 to non-delusional
(e) non-delusional to G.3.1
Figure 4: Breakdown of E.3.2.1 Errors of Skipper on SSM: each mean curve and CI (95%) bar are over $20$ seed runs. In each subfigure, the evolution of $L_{1}$ errors against the ground truth is presented, with each data point computed based on the evolving policy.
B.1.2 Breakdown of Task Performance
In Fig. 5, we present the evolution of Skipper variants’ performance on the training tasks as well as the OOD evaluation tasks throughout the training process. Note that Fig. 3 h) is an aggregation of performances in Fig. 5 b-e).
From the performance advantages of the hybrid variants (in both training and evaluation tasks), we can see that learning to address delusions during training brings better understanding for novel situations posed in OOD tasks.
(a) training, $\delta=0.4$
(b) OOD eval., $\delta=0.25$
(c) OOD eval., $\delta=0.35$
(d) OOD eval., $\delta=0.45$
(e) OOD eval., $\delta=0.55$
Figure 5: Evolution of Performances of Skipper Variants on SSM: each mean curve and CI (95%) bar are over $20$ seed runs.
B.2 Skipper on RDS (Set 2 / 4)
The second environment we employ is RandDistShift, abbreviated as RDS. RDS was originally proposed in Zhao et al. (2021) as a variant of the counterparts in the MiniGrid Baby-AI platform (Chevalier-Boisvert et al., 2018a; b; Hui et al., 2020) and then later used as the experimental backbone in Zhao et al. (2024). We can view RDS as a simpler version of SSM, where everything happens in semantic class $\langle 1,1\rangle$, i.e., agents always spawn with the sword and the shield in hand, thus can acquire the terminal sparse reward by simply navigating to the goal. RDS instances thus have smaller state spaces than its SSM counterparts. The most important difference, in the views of this paper, is that RDS removed the challenges introduced by temporary unreachabilities. This means that G.3.1.1 and E.3.2.2 are no longer relevant, shifting the dominance towards G.3.1 + E.3.2.1 combination. Using RDS not only showcase the performance of the proposed strategies on a controlled environment with G.3.1 + E.3.2.1 dominance, contrasting the G.3.1.1 + E.3.2.2 dominance of SSM, it also can be used to validate the performance of our adapted agents, on an environment where previous benchmarks exist.
We present Skipper’s evaluative curves in Fig. 6.
(a) G.3.1 Candidate Ratio
(b) E.3.2.1 Estimation Errors
(c) G.3.1 + E.3.2.1 Behavior Ratio
(d) Non-Delusional Estim. Errors
(e) Aggregated OOD Performance
Figure 6: Skipper on RDS: each mean curve and CI (95%) bar are over $20$ seed runs. Subfigures b-e) share the same legends. Subfigures a-c) focus on G.3.1 & E.3.2.1 related delusions: a) Evolving ratio of G.3.1 among targets at each candidate selection throughout the training process is presented; b) E.3.2.1 delusions in terms of $L_{1}$ error in estimated distance is visualized, throughout the training process. c) The curves represent the frequencies of choosing G.3.1 targets whenever a selection of targets is initiated; d) The final estimation accuracies of non-delusional source-target pairs after training completed, across a spectrum of ground truth distances. In this figure, both distances (estimation and ground truth) are conditioned on the final version of the evolving policies; The state structure of RDS does not permit G.3.1.1 targets and the corresponding E.3.2.2 delusions; e) Each data point represents OOD evaluation performance aggregated over $4\times 20$ newly generated tasks, with mean difficulty matching the training tasks.
From Fig. 6 e), we can see that, probably because of the lack of dominant G.3.1.1 + E.3.2.2 cases, the OOD performance of even the most basic “episode” variant is high, despite the hybrid variants perform even better. “F-(E+G)”, i.e. the hybrid with the most investment in “generate” (aiming at E.3.2.1), performs the best both in terms of E.3.2.1 delusion suppression (b)), and OOD generalization (e)), as expected. In RDS, the short-distance non-delusional estimation accuracy as well as the OOD performance of “F-P” are not as bad as in SSM. This is possibly due to the fact that RDS has much smaller state spaces, where “episode” and “pertask” produce more similar results (than in large state spaces of SSM).
B.2.1 Breakdown of E.3.2.1 Errors
Similarly, for RDS, we provide a breakdown of E.3.2.1 errors for the Skipper variants in Fig. 7. Similar observations can be made as for SSM.
(a) Overall
(b) from G.3.1 only
(c) to G.3.1 only
(d) G.3.1 to non-delusional
(e) non-delusional to G.3.1
Figure 7: Breakdown of E.3.2.1 Errors of Skipper on RDS: each mean curve and CI (95%) bar are over $20$ seed runs. In each subfigure, the evolution of $L_{1}$ errors against the ground truth is presented, with each data point computed based on the evolving policy.
B.2.2 Breakdown of Task Performance
In Fig. 8, we present the evolution of Skipper variants’ performance on the training tasks as well as the OOD evaluation tasks throughout the training process. Note that Fig. 6 e) is an aggregation of performances in Fig. 8 b-e).
(a) training, $\delta=0.4$
(b) OOD evaluation, $\delta=0.25$
(c) OOD evaluation, $\delta=0.35$
(d) OOD evaluation, $\delta=0.45$
(e) OOD evaluation, $\delta=0.55$
Figure 8: Evolution of Performances of Skipper Variants on RDS: each mean curve and CI (95%) bar are over $20$ seed runs.
B.3 LEAP on SSM (Set 3 / 4)
The third set of experiments, similar to the previous two, is a comparative study for LEAP variants’ performance on SSM. LEAP is different from Skipper, as its decision-time planning process constructs a singular sequence of subgoals leading to the task goal. Due to a lack of backup subgoals, even if one among them is problematic, the whole resulting plan would be delusional, making LEAP much more prone to failures compared to Skipper, where candidate targets can still be reused if deviation from the original plan occurred.
SSM has a relatively large state space that requires more intermediate subgoals for LEAP’s plans. However, an increment of the number of subgoals also dramatically increases the frequencies of delusional plans, damaging the agents’ performance. Because of this, our experimental results of LEAP on SSM with size $12\times 12$ became difficult to analyze because of the rampant failures. We chose instead to present the results on SSM with size $8\times 8$ here.
(a) G.3.1 Ratio in Planned Sequence
(b) G.3.1.1 Ratio in Planned Sequence
(c) Delusional Plan Ratio
(d) Non-Delusional Estimator Errors
(e) Aggregated OOD Performance
Figure 9: LEAP on SSM: each mean curve and CI (95%) bar are over $20$ seed runs. a) Ratio of G.3.1 subgoals among the evolved sequence; b) Ratio of G.3.1.1 subgoals in the evolved sequence; c) Ratio of evolved sequences containing at least one G.3.1 or G.3.1.1 target; d) The final estimation accuracies of non-delusional source-target pairs after training completed, across a spectrum of ground truth distances. In this figure, both distances (estimation and ground truth) are conditioned on the final version of the evolving policies; e) Each data point represents OOD evaluation performance aggregated over $4\times 20$ newly generated tasks, with mean difficulty matching the training tasks.
For LEAP, we use some different metrics to analyze the effectiveness of the proposed strategies in addressing delusions. This is because, if LEAP’s estimator successfully addressed delusions and learned not to favor the problematic targets (G.3.1 and G.3.1.1), then they will not be selected in the evolved elitest sequence of subgoals. This makes it inconvenient for us to use the distance error in the delusional source-target pairs during decision-time as a metric to analyze the reduction of delusional estimates, because of their growing scarcity.
As we can see from Fig. 9, similar arguments about the effectiveness of the proposed hybrid strategies can be made, to those with Skipper. The hybrids with more investment in addressing E.3.2.1, i.e., “F-(E+G)” and “F-(E+P+G)”, exhibit the lowest E.3.2.1 errors (a)). Similarly, “F-(E+P)” and “F-(E+P+G)” achieve the lowest E.3.2.2 errors (b)). In e), we see that the $3$ hybrid variants achieve better OOD performance than the baseline “F-E”. Specifically, “F-(E+G)” achieved the best performance. This is likely because that it induced the highest sample efficiency in terms of learning the estimations between non-delusional subgoals, as shown in d). Assistive strategies such as “generate” and “pertask” do not only induce problematic targets, but also non-delusional ones that can shift the training distribution towards higher sample efficiencies in the traditional sense.
B.3.1 Breakdown of Task Performance
In Fig. 10, we present the evolution of LEAP variants’ performance on the training tasks as well as the OOD evaluation tasks throughout the training process. Note that Fig. 9 e) is an aggregation of performances in Fig. 10 b-e).
(a) training, $\delta=0.4$
(b) OOD evaluation, $\delta=0.25$
(c) OOD evaluation, $\delta=0.35$
(d) OOD evaluation, $\delta=0.45$
(e) OOD evaluation, $\delta=0.55$
Figure 10: Evolution of Performances of LEAP Variants on SSM: each mean curve and CI (95%) bar are over $20$ seed runs.
B.4 LEAP on RDS (Set 4 / 4)
The last set of experiments focus on LEAP’s performance on RDS. Similarly, we present the evaluative metrics in Fig. 11.
(a) G.3.1 Ratio in Planned Sequence
(b) Delusional Plan Ratio
(c) Non-Delusional Estimator Errors
(d) Aggregated OOD Performance
Figure 11: LEAP on RDS: each mean curve and CI (95%) bar are over $20$ seed runs. a) Ratio of G.3.1 subgoals among the evolved sequence; b) Ratio of evolved sequences containing at least one G.3.1 target; c) The final estimation accuracies of non-delusional source-target pairs after training completed, across a spectrum of ground truth distances. In this figure, both distances (estimation and ground truth) are conditioned on the final version of the evolving policies; d) Each data point represents OOD evaluation performance aggregated over $4\times 20$ newly generated tasks, with mean difficulty matching the training tasks.
The conclusions are similar, despite that the OOD performance gain by addressing delusions is significantly higher than in SSM.
B.4.1 Breakdown of Task Performance
In Fig. 12, we present the evolution of LEAP variants’ performance on the training tasks as well as the OOD evaluation tasks throughout the training process. Note that Fig. 11 d) is an aggregation of performances in Fig. 12 b-e).
(a) training, $\delta=0.4$
(b) OOD evaluation, $\delta=0.25$
(c) OOD evaluation, $\delta=0.35$
(d) OOD evaluation, $\delta=0.45$
(e) OOD evaluation, $\delta=0.55$
Figure 12: Evolution of Performances of LEAP Variants on SSM: each mean curve and CI (95%) bar are over $20$ seed runs.
B.5 Discussions & More Details of “generate” & “pertask
Figure 13: Representative Atom Hindsight Relabeling Strategies & Newly Proposed Ones: The first two strategies, “future” and “episode”, are widely used as they create relabeled transitions that help estimators efficiently handle non-delusional planning. The last two, “generate” and “pertask”, are effective at addressing delusions, making them useful in specific scenarios. Atomic hindsight strategies from the first group can serve as backbones for mixture strategies, complemented by the second group to address delusions.
B.5.1 Implementation pf “pertask
pertask” takes the advantage for the fact that training is done on limited number of fixed task instances. We give each task a unique task identifier. At relabeling time, “pertask” samples observations among all the transitions marked with the same identifier as the current training task instance. This can be trivially implemented with individual auxiliary ERs that store only the experienced states with memory-efficient pointers to the buffered $x_{t}$’s in the main HER.
B.5.2 Discussions
generate” not only creates targets with G.3.1 targets, but also generate valid targets that should resemble the distribution it was trained on. Thus, it is not clear if mixing in data augmented by “generate” would result in lower sample efficiency in the estimation cases involving valid targets. Take SSM as an example, “generate” seemed to have detrimental effect to non-delusional cases when applied to Skipper, while it greatly boosted accuracies for LEAP overall.
In some experiments, “pertask” demonstrated clear effectiveness in addressing E.3.2.1 as well, despite that it was not designed to. This is likely because of some generalization effects of the estimator, which were trained with additional data that boosted data diversity.
In some environments, we expect that “pertask” could also be used (for mixtures of the generator) to learn to generate longer-distance targets from the current states if the generator has trouble doing so with “future”, with the accompanied risks of lower efficiency and G.3.1.1 hallucinations.
Appendix C Implementation Details for Experiments
C.1 Skipper
Our adaptation of Skipper over the original implementation [^2] https://github.com/mila-iqia/Skipper in Zhao et al. (2024) is minimal. We have additionally added two simple vertex pruning procedures before the vertex pruning based on $k$-medoids. These two procedures include: 1) prune vertices that are duplicated, and 2) prune vertices that cannot be reached from the current state with the estimated connectivity.
We implemented a version of generator that can reliably handle both RDS and SSM with the same architecture. Please consult models.py in the submitted source code for its detailed architecture.
For SSM instances, since their state spaces are much bigger than those of RDS, we ask that Skipper generate twice the number of candidate targets (both before and after pruning) for the proxy problems.
All other architectures and hyperparameters are identical to the original implementation.
For better adaptability during evaluation and faster training, Skipper variants in this paper keeps the constructed proxy problem for the whole episode during training and replanning only triggers a re-selection, while during evaluation, the proxy problems are always erased and re-constructed.
The quality of our adaptation of the original implementation can be assured by the fact the “F-E” variant’s performance matches the original on RDS.
C.2 LEAP
C.2.1 Adaptation for Discrete Action Spaces
LEAP’s training involves two pretraining stages, that are, generator pretraining and distance estimator training.
We improved upon the adopted discrete-action space compatible implementation of LEAP (Nasiriany et al., 2019) from Zhao et al. (2024). We gave LEAP additional flexibility to use fewer subgoals along the way to the task goal if necessary. Also, we improved upon the Cross-Entropy Method (CEM), such that elite sequences would be kept intact in the next population during the optimization process. We increased the base population size of each generation to $512$ and lengthened the number of iterations to $10$.
For RDS $12\times 12$ and SSM $8\times 8$, at most $3$ subgoals are used in each planned trajectory. We find that employing more subgoals greatly increases the burden of evolutionary computations and lower the quality of the evolved subgoal sequences, leading to bad performance that cannot be effectively analyzed.
We used the same generator architecture and hyperparameters as in Skipper. All other architectures and hyperparameters remain unchanged.
Similarly for LEAP, for better adaptability during evaluation, the planned sequences of subgoals are always reconstructed whenever planning is triggered. While in training, the sequence is reused and only a subgoal selection is conducted.
The quality of our adaptation of the original implementation can be assured by the fact the “F-E” variant’s performance matches the original on RDS.